OpenStudy (anonymous):

Using the quadratic formula, Solve this equation h^2-2h=8 .. So x = ?

OpenStudy (solomonzelman):

\(\large\color{slate}{ h^2-2h=8 }\) add 1 to both sides, \(\large\color{slate}{ h^2-2h\color{red}{+1}=8\color{red}{+1} }\) you get: \(\large\color{slate}{ h^2-2h+1=8+1 }\) NOW, complete the square.

OpenStudy (solomonzelman):

(And you know what 8+1 is)

OpenStudy (anonymous):

9 !

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Dont you need to square root, 3 is not the right answer.. so confused

OpenStudy (misty1212):

if you really want to use the quadratic formula, rather than completing the square (which would be easier in this case) then the first thing you need to do is write \[h^2-2h-8=0\] then use \[h=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=1,b=-2,c=-8\]

OpenStudy (solomonzelman):

you can factor it

OpenStudy (misty1212):

it would be easier to complete the square, still easier to factor \[h^2-2h-8=(h+2)(h-4)\]

OpenStudy (solomonzelman):

completing the square was the first that came on my mind.